OpenStudy (anonymous):

An object is dragged 5m along the ground by a 47 N force acting at 50 degrees to the ground. It is then dragged 8 m up a ramp inclined at 30 degrees by the same exact force. Find the total work done.

OpenStudy (anonymous):

wouldn't it be \[Fd _{1} \cos \theta _{1}+Fd _{2}\cos \theta _{2}\]

OpenStudy (anonymous):

where F is force and d is distance in meters

OpenStudy (anonymous):

yeah but I'd like to know how exactly I should draw the diagram

OpenStudy (e.cociuba):

Use this formula... Work=Force*distance Where work is expressed in joules, force in newtons, and distance in meters.

OpenStudy (anonymous):

idk that's a good question because are you still exerting the forrce at a 50 degree incline while on the 30 degree incline on the ramp?

OpenStudy (anonymous):

I'm trying to help you solve it while in the midst of learning it too lol I'm in Physics right now as well

OpenStudy (anonymous):

Ah tbh I really don't know since it doesn't say so in the question unfortunately :( it just asks for how much work is done in TOTAL

OpenStudy (anonymous):

hmmmm. idk how many people are good with physics in this section but there's gotta b someone

OpenStudy (e.cociuba):

Haha oops I thought It was a physics question. Srry @TheSweetLegend

OpenStudy (e.cociuba):

Isn't it??

OpenStudy (anonymous):

@ZeHanz

OpenStudy (anonymous):

it's ok @e.cociuba! :) actually that was a formula my vectors teacher recommended the class using but I was curious as to why the angles were given.

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

apparently this is what happens when you decide to not take physics :(

OpenStudy (zehanz):

First part: \(W=F \cdot s=47\cos50⁰\cdot5\approx151.06~ J\)

OpenStudy (e.cociuba):

Ohh ok I see.. lol @TheSweetLegend

OpenStudy (anonymous):

Thanks @ZeHanz! :D so the 2nd part I do the same thing except for the ramp? and then I add them up to get the final answer right?

OpenStudy (zehanz):

Yes, although I've not worked out yet how to calculate that second part. There is a distance covered of 8 m, by a force of 47cos(20⁰). But: the object is also moved up by 4 m against gravity, so I'd say that also involves some work...

OpenStudy (anonymous):

oooohhh... so how should I solve that part?

OpenStudy (anonymous):

and by 20 you mean 30 degrees from the ramp right?

OpenStudy (zehanz):

the 20 degrees is with respect to the ramp (50-30=20). Maybe it is better to calculate the work done by the horizontal movement and by the vertical movement separately. Horizontal: \(4\sqrt{3}~m\cdot 47\cos50⁰~N \approx209.31~J\) Vertical \(4 \cdot47\sin50⁰~N \approx 144.02~J \) Here, I have basically ignored the ramp and just looked at the point where the object arrives. Now I'm not sure if I can just add the 209.31 and 144.02, or that I have to use the pythagorean theorem...